OpenStudy (anonymous):

A rectangular field is to be enclosed on four sides with a fence. Fencing cost $5/ft on two sides. Other sides cost $8/ft. Find the dim of the field with an A=830ft^2 that is the cheapest to enclose?

OpenStudy (anonymous):

dim = dimension ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

l=830/w thats about as far as ive gotten

OpenStudy (anonymous):

w=28.8?

OpenStudy (anonymous):

we need to find the prime factors of 830 in order to find combination of 2 min possible multiples of 830 because then only we will find cheapest endorsement we can do it like this let's take length=l and width =w so 830=l*w as u said l=830/w now cost of fencing=8*w +5*l put l=830/w here so cost=8*w+(5*830)/w now differntiate it with respect to w and since we want to minimize it equate this expression to 0